When x number of units are sold, the price of each unit(in dollars)is given by p=-x/2+75. find the unit price when the following quantities are sold:2,7,9,11.
Answer by checkley71(8403)

P(2)=-2/2+75 P(2)=-1+75 P(2)= 74
P(7)=-7/2+75 P(7)=-3.50+75 P(4)=71.50
P(9)=-9/2+75 P(9)=-4.50+75 P(9)=70.50
P(11)=-11/2+75 P(11)=-5.5+75 P(11)=69.50